OpenStudy (anonymous):

A student lifts a 20.0 kg box 1.5 m straight up at a constant speed. Find the work done by each force that acts on the box.

OpenStudy (anonymous):

I tried it and got (20)(9.8)(1.5) for Applied Force- would it just be negative for Fg?

OpenStudy (roadjester):

\(\large W=F\Delta x cos\theta\)

OpenStudy (roadjester):

I'm actually confusing myself... The applied force is upward due to lifting upward but gravitational acceleration is downward... @theEric

OpenStudy (roadjester):

WAIT, acceleration is 0! there is no work upward since it's constant velocity... I think...

OpenStudy (theeric):

There is no work done on the box, but that's because the work done by the lifter and the work done by gravity are equal in magnitude and opposite direction. So they cancel! \(W=Fd\cos\theta\) where \(\theta\) is the angle between the force and displacement.

OpenStudy (anonymous):

So I was right then? The work done by Fg is just negative work done by Fa?

OpenStudy (theeric):

@greenglasses You are right!

OpenStudy (theeric):

Yep!

OpenStudy (roadjester):

Thanks @theEric I was confusing the heck out of myself

OpenStudy (anonymous):

Thanks, my teacher didn't leave us any answers so I had no way to check.

OpenStudy (theeric):

Haha, no problem. I dearly hope I'm not confused! :)

OpenStudy (theeric):

Cool! I'm glad OpenStudy is here.